# KazSAnDRA: Kazakh Sentiment Analysis Dataset of Reviews and Attitudes

### v1

@@ Abstract
KazSAnDRA comprises an extensive collection of 180,064 reviews obtained from various sources and includes numerical ratings ranging from 1 to 5, providing a quantitative representation of customer attitudes. The study also pursued the automation of Kazakh sentiment classification through the development and evaluation of four machine learning models trained for both polarity classification and score classification. Experimental analysis included evaluation of the results considering both balanced and imbalanced scenarios. The most successful model attained an F1-score of 0.81 for polarity classification and 0.39 for score classification on the test sets. The dataset and fine-tuned models are open access and available for download under the Creative Commons Attribution 4.0 International License (CC BY 4.0) through our GitHub repository.
@@ 1. Introduction
We also developed and evaluated four machine learning models to automate the classification of Kazakh sentiments. The highest F1-score on the test sets was 0.81 for polarity classification and 0.39 for score classification.
@@ 3.8. Dataset Organisation
| Model | Polarity Classification | Score Classification | |||||||||||||||||||||||||||
| Balanced (ROS) | Balanced (RUS) | Imbalanced | Balanced (ROS) | Balanced (RUS) | Imbalanced | ||||||||||||||||||||||||
| A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | A | P | R | F1 | ||||||
| mBERT | 0.84 | 0.74 | 0.83 | 0.77 | 0.85 | 0.76 | 0.82 | 0.78 | 0.89 | 0.82 | 0.79 | 0.80 | 0.67 | 0.34 | 0.36 | 0.35 | 0.63 | 0.35 | 0.39 | 0.36 | 0.77 | 0.44 | 0.36 | 0.37 | |||||
| XLM-R | 0.86 | 0.76 | 0.83 | 0.79 | 0.85 | 0.75 | 0.83 | 0.78 | 0.89 | 0.81 | 0.81 | 0.81 | 0.58 | 0.36 | 0.42 | 0.36 | 0.66 | 0.36 | 0.41 | 0.37 | 0.77 | 0.42 | 0.37 | 0.39 | |||||
| RemBERT | 0.88 | 0.79 | 0.82 | 0.81 | 0.87 | 0.78 | 0.82 | 0.80 | 0.89 | 0.81 | 0.82 | 0.81 | 0.73 | 0.37 | 0.36 | 0.36 | 0.62 | 0.35 | 0.40 | 0.35 | 0.76 | 0.41 | 0.38 | 0.39 | |||||
| mBART-50 | 0.87 | 0.77 | 0.79 | 0.78 | 0.81 | 0.72 | 0.81 | 0.74 | 0.89 | 0.82 | 0.78 | 0.80 | 0.74 | 0.36 | 0.34 | 0.35 | 0.55 | 0.36 | 0.41 | 0.34 | 0.77 | 0.42 | 0.37 | 0.38 | |||||
@@ 4.2. Experimental Setup
All four models were fine-tuned using both the balanced and imbalanced training sets, while the hyperparameters were refined using the validation set. The final and most optimal models were evaluated on the test sets. The fine-tuning of the models was executed on a single A100 GPU hosted on an NVIDIA DGX A100 machine. The initial learning rate was set at ; the weight decay rate was set at . Early stopping was employed, executed when the F1-score exhibited no improvement for three consecutive epochs. We set the batch size to 32 (mBERT, XLM-R, RemBERT) or 16 (mBART-50) and applied 800 warm-up steps.
@@ 4.3. Performance Metrics
Several conventional metrics were used to evaluate the performance of the models, including accuracy (A), precision (P), recall (R), and F1-score (F1). Given the imbalanced nature of the dataset, where all classes carry equal importance, we opted for macro-averaging, calculated from the arithmetic (i.e., unweighted) mean of all F1-scores per class, and thus ensuring equal treatment of all classes during the evaluation, resulting in a stronger penalty if the model performs worse on minority classes (Jurafsky and Martin, 2009; Yang, 2001).
@@ 4.4. Experiment Results
Table 10 presents the performance of the four models on KazSAnDRA test sets for the PC and SC tasks. XLM-R and RemBERT consistently outperformed mBERT and mBART-50 across various training scenarios. The highest F1-scores of 0.81 (PC) and 0.39 (SC) were achieved by both XLM-R and RemBERT in the imbalanced training scenario. In five out of six training scenarios, RemBERT achieved the highest F1-scores, while XLM-R led in four. Table 11 presents data on the number of epochs required to train models for the PC and SC tasks, considering both balanced (ROS, RUS) and imbalanced (IB) training data scenarios. Tables 12–14 summarise the performance of RemBERT in the imbalanced training context, with a detailed analysis following in the subsequent section.
| Model | PC | SC | |||||
| ROS | RUS | IB | ROS | RUS | IB | ||
| mBERT | 4 | 7 | 6 | 8 | 10 | 11 | |
| XLM-R | 5 | 7 | 5 | 4 | 9 | 16 | |
| RemBERT | 4 | 5 | 5 | 6 | 6 | 9 | |
| mBART-50 | 5 | 7 | 5 | 8 | 7 | 5 | |
@@ 5. Discussion
| Domain | PC | SC | |||||||
| A | P | R | F1 | A | P | R | F1 | ||
| Appstore | 0.87 | 0.80 | 0.81 | 0.80 | 0.74 | 0.41 | 0.37 | 0.38 | |
| Bookstore | 0.86 | 0.75 | 0.80 | 0.77 | 0.73 | 0.34 | 0.32 | 0.32 | |
| Mapping | 0.92 | 0.84 | 0.88 | 0.86 | 0.80 | 0.42 | 0.41 | 0.41 | |
| Market | 0.97 | 0.84 | 0.91 | 0.87 | 0.82 | 0.43 | 0.41 | 0.42 | |
@@ 6. Conclusion
The aim of this study was to create an extensive dataset for Kazakh sentiment analysis. The result is KazSAnDRA, the first and largest publicly available dataset for Kazakh. Comprising 180,064 reviews from four domains, KazSAnDRA includes numerical ratings from 1 to 5 that quantitatively represent customers’ attitudes. To automate Kazakh sentiment classification, we developed and evaluated four machine learning models for both polarity and score classification. The experimental analysis involved examining the results obtained with both balanced and imbalanced training data. The most successful model achieved an F1-score of 0.81 for the polarity classification task and 0.39 for the score classification task on the test sets. In the future, we plan to improve KazSAnDRA by addressing spelling errors and effectively handling code-switching phenomena. These improvements will facilitate the use of advanced data augmentation techniques to cope with data imbalance challenges.

### v2

@@ Abstract
KazSAnDRA comprises an extensive collection of 180,064 reviews obtained from various sources and includes numerical ratings ranging from 1 to 5, providing a quantitative representation of customer attitudes. The study also pursued the automation of Kazakh sentiment classification through the development and evaluation of four machine learning models trained for both polarity classification and score classification. Experimental analysis included evaluation of the results considering both balanced and imbalanced scenarios. The most successful model attained an F-score of 0.81 for polarity classification and 0.39 for score classification on the test sets. The dataset and fine-tuned models are open access and available for download under the Creative Commons Attribution 4.0 International License (CC BY 4.0) through our GitHub repository.
@@ 1. Introduction
We also developed and evaluated four machine learning models to automate the classification of Kazakh sentiments. The highest F-score on the test sets was 0.81 for polarity classification and 0.39 for score classification.
@@ 3.8. Dataset Organisation
| Model | Polarity Classification | Score Classification | ||||||||||||||||||||||
| Balanced (ROS) | Balanced (RUS) | Imbalanced | Balanced (ROS) | Balanced (RUS) | Imbalanced | |||||||||||||||||||
| A | P | R | F | A | P | R | F | A | P | R | F | A | P | R | F | A | P | R | F | A | P | R | F | |
| mBERT | 0.84 | 0.74 | 0.83 | 0.77 | 0.85 | 0.76 | 0.82 | 0.78 | 0.89 | 0.82 | 0.79 | 0.80 | 0.67 | 0.34 | 0.36 | 0.35 | 0.63 | 0.35 | 0.39 | 0.36 | 0.77 | 0.44 | 0.36 | 0.37 |
| XLM-R | 0.86 | 0.76 | 0.83 | 0.79 | 0.85 | 0.75 | 0.83 | 0.78 | 0.89 | 0.81 | 0.81 | 0.81 | 0.58 | 0.36 | 0.42 | 0.36 | 0.66 | 0.36 | 0.41 | 0.37 | 0.77 | 0.42 | 0.37 | 0.39 |
| RemBERT | 0.88 | 0.79 | 0.82 | 0.81 | 0.87 | 0.78 | 0.82 | 0.80 | 0.89 | 0.81 | 0.82 | 0.81 | 0.73 | 0.37 | 0.36 | 0.36 | 0.62 | 0.35 | 0.40 | 0.35 | 0.76 | 0.41 | 0.38 | 0.39 |
| mBART-50 | 0.87 | 0.77 | 0.79 | 0.78 | 0.81 | 0.72 | 0.81 | 0.74 | 0.89 | 0.82 | 0.78 | 0.80 | 0.74 | 0.36 | 0.34 | 0.35 | 0.55 | 0.36 | 0.41 | 0.34 | 0.77 | 0.42 | 0.37 | 0.38 |
@@ 4.2. Experimental Setup
All four models were fine-tuned using both the balanced and imbalanced training sets, while the hyperparameters were refined using the validation set. The final and most optimal models were evaluated on the test sets. The fine-tuning of the models was executed on a single A100 GPU hosted on an NVIDIA DGX A100 machine. The initial learning rate was set at ; the weight decay rate was set at . Early stopping was employed, executed when the F-score exhibited no improvement for three consecutive epochs. We set the batch size to 32 (mBERT, XLM-R, RemBERT) or 16 (mBART-50) and applied 800 warm-up steps.
@@ 4.3. Performance Metrics
Several conventional metrics were used to evaluate the performance of the models, including accuracy (A), precision (P), recall (R), and F-score (F). Given the imbalanced nature of the dataset, where all classes carry equal importance, we opted for macro-averaging, calculated from the arithmetic (i.e., unweighted) mean of all F-scores per class, and thus ensuring equal treatment of all classes during the evaluation, resulting in a stronger penalty if the model performs worse on minority classes (Jurafsky and Martin, 2009; Yang, 2001).
@@ 4.4. Experiment Results
Table 10 presents the performance of the four models on KazSAnDRA test sets for the PC and SC tasks. XLM-R and RemBERT consistently outperformed mBERT and mBART-50 across various training scenarios. The highest F-scores of 0.81 (PC) and 0.39 (SC) were achieved by both XLM-R and RemBERT in the imbalanced training scenario. In five out of six training scenarios, RemBERT achieved the highest F-scores, while XLM-R led in four. Table 11 presents data on the number of epochs required to train models for the PC and SC tasks, considering both balanced (ROS, RUS) and imbalanced (IB) training data scenarios. Tables 12–14 summarise the performance of RemBERT in the imbalanced training context, with a detailed analysis following in the subsequent section.
| Model | PC | SC | ||||
| ROS | RUS | IB | ROS | RUS | IB | |
| mBERT | 4 | 7 | 6 | 8 | 10 | 11 |
| XLM-R | 5 | 7 | 5 | 4 | 9 | 16 |
| RemBERT | 4 | 5 | 5 | 6 | 6 | 9 |
| mBART-50 | 5 | 7 | 5 | 8 | 7 | 5 |
@@ 5. Discussion
| Domain | PC | SC | ||||||
| A | P | R | F | A | P | R | F | |
| Appstore | 0.87 | 0.80 | 0.81 | 0.80 | 0.74 | 0.41 | 0.37 | 0.38 |
| Bookstore | 0.86 | 0.75 | 0.80 | 0.77 | 0.73 | 0.34 | 0.32 | 0.32 |
| Mapping | 0.92 | 0.84 | 0.88 | 0.86 | 0.80 | 0.42 | 0.41 | 0.41 |
| Market | 0.97 | 0.84 | 0.91 | 0.87 | 0.82 | 0.43 | 0.41 | 0.42 |
@@ 6. Conclusion
The aim of this study was to create an extensive dataset for Kazakh sentiment analysis. The result is KazSAnDRA, the first and largest publicly available dataset for Kazakh. Comprising 180,064 reviews from four domains, KazSAnDRA includes numerical ratings from 1 to 5 that quantitatively represent customers’ attitudes. To automate Kazakh sentiment classification, we developed and evaluated four machine learning models for both polarity and score classification. The experimental analysis involved examining the results obtained with both balanced and imbalanced training data. The most successful model achieved an F-score of 0.81 for the polarity classification task and 0.39 for the score classification task on the test sets. In the future, we plan to improve KazSAnDRA by addressing spelling errors and effectively handling code-switching phenomena. These improvements will facilitate the use of advanced data augmentation techniques to cope with data imbalance challenges.